Vector A and B have a Scalar product of -10 and their vector product has a magnitude of 8.5. What is the angle between these two vectors?
1
Expert's answer
2019-07-24T15:49:14-0400

The scalar product of two vectors

"({\vec A}\cdot{\vec B})=AB\cos\theta"

The magnitude of vector product of two vectors

"|({\vec A}\times{\vec B})|=AB\sin\theta"

So

"\tan\theta=\frac{|({\vec A}\times{\vec B})|}{({\vec A}\cdot{\vec B})}"

"\tan\theta=\frac{8.5}{-10}=-0.85"

Finally, the angle between vectors A and B

"\theta=\tan^{-1}(-0.85)=139.6^{\circ}"